# Kang Chen♣,♡,∗, Zhihao Liu♢,♡,∗, Tonghe Zhang♯,⊳,∗, Zhen Guo, Si Xu, Hao Lin, Hongzhi Zang, & Xiang Li, Quanlu Zhang, Zhaofei Yu, Guoliang Fan, Tiejun Huang, Yu Wang♠,†, Chao Yu♠,♡,†

### v2

@@ 4.2 Flow-SDE
Inspired by Flow-GRPO (liu2025flowgrpo), we enhance stochastic exploration by converting the denoising process from ODE into an SDE formulation. We further construct a two-layer MDP to couple the denoising process with the policy-environment interaction following DPPO (ren2024dppo), while leveraging the hybrid ODE-SDE sampling technique to accelerate the training process.
@@ 4.2.1 Stochasticity Injection
In Flow-SDE, we convert the deterministic ODE into an equivalent SDE that preserves the marginal probability density of the generated actions, as shown in Fig.˜3.
The deterministic ODE sampling trajectory of the flow matching, especially the Rectified Flow (liu2022recflow), is described by the forward Euler method:
@@
Building on the connection between the probability flow ODE and SDE (song2020score), we can transform the deterministic ODE in Eq.˜6 into an equivalent SDE, with a drift term that corrects the original velocity and a diffusion term that introduces noise:
@@ 4.2.2 MDP Formulation
While Flow-Noise substitutes the joint log-likelihood of the entire denoising sequence for the likelihood of the final executed action, we couple the denoising process of the flow matching with environmental interaction in Flow-SDE. Specifically, we embed the inner MDP defined during the denoising process into the high-level, outer-loop MDP with the environment $\mathcal{M}_{\text{ENV}}$ in Sec.˜3.1, formulating a two-layer MDP as shown in Fig.˜2, with components defined with respect to the environment time $t$ and denoising time $\tau$.
@@ 4.2.3 Hybrid ODE-SDE Sampling
In the formulated two-layer MDP framework, the effective trajectory length is the product of the environment interaction steps and the number of flow matching denoising steps. While this formulation enables RL training for flow-based VLAs, it significantly extends the MDP horizon compared to non-iterative VLA methods, which substantially increases both the training difficulty and the computational time required for optimization.
To this end, we adopt the mixed ODE-SDE rollout strategy, drawing inspiration from the text-to-image generation methods such as Mix-GRPO (li2025mixgrpo) and TempFlow-GRPO (he2025tempflow). Specifically, during the denoising process, a single step is randomly sampled as a stochastic SDE transition governed by $p(\mathbf{A}^{\tau+\delta}|\mathbf{A}^{\tau})\sim\mathcal{N}(\mu_{\tau},\Sigma_{\tau})$, while the remaining steps follow deterministic ODE transitions defined by the update rule $\mathbf{A}^{\tau+\delta}=\mathbf{A}^{\tau}+\mathbf{v}^{\tau}\cdot\delta$.
Under this formulation, we treat the deterministic ODE transition between states as an environment-level wrapper and revise the state transition function of the previous two-layer MDP. Specifically, at each environment step $t$, a denoising time $\tau_{t}$ is randomly selected for the policy’s stochastic injection. The policy $\pi$ acts on this state $\bar{s}_{t}^{\tau_{t}}=(\mathbf{o}_{t},\mathbf{A}_{t}^{\tau_{t}})$, sampling the action $\mathbf{A}_{t}^{\tau_{t}+\delta}$ according to Eq.˜10. The environment wrappers then execute all subsequent deterministic steps, ultimately transitioning to the next observation $\mathbf{o}_{t+1}$ and the next action state $\bar{s}_{t+1}^{\tau_{t+1}}=(\mathbf{o}_{t+1},\mathbf{A}_{t+1}^{\tau_{t+1}})$ at a newly sampled time $\tau_{t+1}$. During this process, the state input and action output of the policy remain consistent with the previous two-layer MDP formulation, thus ensuring theoretical consistency.
@@ 4.3.1 Algorithm
$\pi$-series models (black2024pi_0; intelligence2025pi05) adopt a chunk-based approach for action generation. Specifically, the policy outputs an entire sequence of $H$ future actions $\mathbf{A}_{t}=[a_{t,0},...,a_{t,H-1}]$ in response to each observation. In this approach, we treat the entire sequence as a single macro-step and define its corresponding reward $R_{t}=\sum_{j=0}^{H-1}r_{t,j}$ as the sum of the per-step rewards $r_{t,j}$, referred to as the chunk-level formulation in RLinf-VLA (zang2025rlinf).
To effectively guide policy updates, PPO employs Generalized Advantage Estimation (GAE) (gae) to compute a low-variance estimate of the advantage, estimated as:
@@
$\rho_{t}(\theta)=\frac{\pi_{\theta_{\text{new}}}(a_{t}|s_{t})}{\pi_{\theta_{\text{old}}}(a_{t}|s_{t})}\quad\text{or}\quad\rho_{t}(\theta)=\frac{\pi_{\theta_{\text{new}}}(\bar{a}_{t}^{\tau}|\bar{s}_{t}^{\tau})}{\pi_{\theta_{\text{old}}}(\bar{a}_{t}^{\tau}|\bar{s}_{t}^{\tau})},$ (15)
for the one-layer and two-layer MDP formulations, respectively.
(a) Critic with the action expert, exemplified by $\pi_{0}$.
(b) Critic with the VLM, exemplified by $\pi_{0.5}$.
Figure 4: Illustration of the two critic placement configurations.
@@ 4.3.2 Critic Design
Following VLA-PPO works (zang2025rlinf; rl4vla), we employ a shared actor-critic architecture for memory-efficient value prediction as shown in Fig.˜4. However, the two flow-based VLAs process the proprioceptive state differently: in $\pi_{0}$, the state is fed into the action expert model, whereas in $\pi_{0.5}$, it is merged with prompt embeddings within the VLM.
@@ 5.1 Setup
Benchmarks. We perform experiments based on LIBERO (liu2023libero), ManiSkill (tao2024maniskill3) and MetaWorld (mclean2025meta) benchmarks.
  • LIBERO (liu2023libero) is built on a CPU-based simulation platform MuJoCo and assesses knowledge transfer in robotic multi-task and lifelong learning across four manipulation task suites: Spatial, Object, Goal, and Long.
  • ManiSkill serves as a high-fidelity, GPU-parallelized simulation platform. Within ManiSkill, we adopt the SIMPLER benchmark (SIMPLER) as our primary testbed. To further evaluate the generalization capability of $\pi_{\texttt{RL}}$, we follow the setup of RL4VLA (rl4vla) and construct 4,352 pick-and-place task combinations as an extended benchmark.
  • MetaWorld (mclean2025meta) is a multi-task evaluation benchmark built upon the MuJoCo simulator. To evaluate performance on a broad spectrum of manipulation skills beyond pick-and-place, we utilize the MT50 task set as our testbed.
Flow-based VLAs. We conduct experiments based on $\pi_{0}$ and $\pi_{0.5}$. $\pi_{0}$ introduces the flow-matching action expert (300M) built upon a pre-trained PaliGemma (3B) to leverage broad semantic knowledge from internet-scale data. $\pi_{0.5}$ further utilizes co-training across heterogeneous data sources (e.g., multi-robot data, web data, and high-level semantic predictions) for broader generalization. In addition to the $\pi$-series models, we also conduct experiments on GR00T (bjorck2025gr00t) in Appendix appendix˜C, validating that our algorithm is applicable to other flow-based VLAs.
@@
Implementation Details. Given that pre-trained models often struggle to generalize to task-specific benchmarks, we initiate our process with SFT on expert demonstrations. For the SFT stage, we fine-tune the entire 3.3B model following the official setting. In the subsequent RL stage, we freeze the VLM parameters and exclusively fine-tune the 300M action expert model, driven by GPU memory efficiency and the findings from RL4VLA that RL contributes more significantly to action generalization. We build the whole framework upon the RLinf (yu2025rlinf) codebase, where we adopt a shared, co-located GPU allocation strategy that places the environment, rollout model, and actor model on the same GPU and executes them serially.
@@
For the model configurations, we adhere to the official setting provided by openpi (black2024pi_0; intelligence2025pi05). In these settings, $\pi_{0}$ utilizes image, language, and proprioceptive states as input, whereas $\pi_{0.5}$ notably omits state information for the LIBERO benchmark [^2] https://github.com/Physical-Intelligence/openpi/issues/687. Following this precedent, we consistently omit the state input for $\pi_{0.5}$ during both SFT and RL phases on LIBERO and ManiSkill. Our experiments are conducted on 8 NVIDIA H100 80GB GPUs, and detailed training hyperparameters are available in Appendix Tabs.˜7 and 8.
@@
5.2 Main Results
5.2.1 LIBERO

### v3

@@ 4.2 Flow-SDE
Inspired by Flow-GRPO (Liu et al., 2025a), we enhance stochastic exploration by converting the denoising process from ODE into an SDE formulation. We further construct a two-layer MDP to couple the denoising process with the policy-environment interaction following DPPO (Ren et al., 2024), while leveraging the hybrid ODE-SDE sampling technique to accelerate the training process.
@@ 4.2.1 Stochasticity Injection
In Flow-SDE, we convert the deterministic ODE into an equivalent SDE that preserves the marginal probability density of the generated actions, as shown in Fig.˜2.
The deterministic ODE sampling trajectory of the flow matching, especially the Rectified Flow (Liu et al., 2022), is described by the forward Euler method:
@@
Building on the connection between the probability flow ODE and SDE (Song et al., 2020), we can transform the deterministic ODE in Eq.˜6 into an equivalent SDE, with a drift term that corrects the original velocity and a diffusion term that introduces noise:
@@ 4.2.2 MDP Formulation
We couple the denoising process of the flow matching with environmental interaction in Flow-SDE. Specifically, we embed the inner MDP defined during the denoising process into the high-level, outer-loop MDP with the environment $\mathcal{M}_{\text{ENV}}$ in Sec.˜3.1, formulating a two-layer MDP as shown in Fig.˜3, with components defined with respect to the environment time $t$ and denoising time $\tau$.
@@
(a) Critic with the action expert, exemplified by $\pi_{0}$.
(b) Critic with the VLM, exemplified by $\pi_{0.5}$.
Figure 4: Illustration of the two critic placement configurations.
@@ 4.2.3 Hybrid ODE-SDE Sampling
The two-layer MDP formulation significantly extends the horizon, increasing training difficulty and computational cost. To mitigate this, we adopt a mixed ODE-SDE rollout strategy (Li et al., 2025b; He et al., 2025). At each step $t$, we randomly sample a denoising time $\tau_{t}$ for the stochastic SDE exploration, while treating all remaining denoising steps as deterministic ODE updates. Specifically, the policy acts on the state $\bar{s}_{t}^{\tau_{t}}=(\mathbf{o}_{t},\mathbf{A}_{t}^{\tau_{t}})$; subsequently, an environment wrapper executes the remaining ODE steps and the environment transition, ultimately yielding the next state $\bar{s}_{t+1}^{\tau_{t+1}}=(\mathbf{o}_{t+1},\mathbf{A}_{t+1}^{\tau_{t+1}})$ at a newly sampled time $\tau_{t+1}$. This formulation effectively shortens the MDP horizon while maintaining theoretical consistency with the original two-layer framework.
@@ 4.3.1 Algorithm
$\pi$-series models (Black et al., 2024; Intelligence et al., 2025) adopt a chunk-based approach for action generation. Specifically, the policy outputs an entire sequence of $H$ future actions $\mathbf{A}_{t}=[a_{t,0},...,a_{t,H-1}]$ in response to each observation. In this approach, we treat the entire sequence as a single macro-step and define its corresponding reward $R_{t}=\sum_{j=0}^{H-1}r_{t,j}$ as the sum of the per-step rewards $r_{t,j}$, referred to as the chunk-level formulation in RLinf-VLA (Zang et al., 2025).
To effectively guide policy updates, PPO employs Generalized Advantage Estimation (GAE) (Schulman et al., 2015) to compute a low-variance estimate of the advantage, estimated as:
@@
$\rho_{t}(\theta)=\frac{\pi_{\theta_{\text{new}}}(a_{t}|s_{t})}{\pi_{\theta_{\text{old}}}(a_{t}|s_{t})}\quad\text{or}\quad\rho_{t}(\theta)=\frac{\pi_{\theta_{\text{new}}}(\bar{a}_{t}^{\tau}|\bar{s}_{t}^{\tau})}{\pi_{\theta_{\text{old}}}(\bar{a}_{t}^{\tau}|\bar{s}_{t}^{\tau})}.$ (15)
@@ 4.3.2 Critic Design
Following VLA-PPO works (Zang et al., 2025; Liu et al., 2025b), we employ a shared actor-critic architecture for memory-efficient value prediction as shown in Fig.˜4. However, the two flow-based VLAs process the proprioceptive state differently: in $\pi_{0}$, the state is fed into the action expert model, whereas in $\pi_{0.5}$, it is merged with prompt embeddings within the VLM.
@@ 5.1 Setup
Benchmarks. We perform experiments on four widely-adopted robot manipulation benchmarks: LIBERO (Liu et al., 2023), ManiSkill (Tao et al., 2024), MetaWorld (McLean et al., 2025) and CALVIN (Mees et al., 2022).
Flow-based VLAs. We conduct our primary experiments based on $\pi_{0}$ and $\pi_{0.5}$ models. Additionally, we conduct experiments on GR00T in Appendix Appendix˜H, which validates that our algorithm can be applied to other flow-based VLAs.
Table 1: Comprehensive ID performance comparison across four benchmarks.
Model Benchmarks
LIBERO ManiSkill MetaWorld CALVIN Avg. $\Delta$ Avg.
$\pi_{0}$ SFT 57.6 38.4 50.8 57.5 51.1
Flow-SDE 96.1 78.8 78.1 61.7 78.7 +27.6
Flow-Noise 97.6 77.8 85.8 59.9 80.3 +29.2
$\pi_{0.5}$ SFT 77.1 40.1 43.8 61.3 55.6
Flow-SDE 97.9 90.9 70.7 87.0 86.6 +31.0
Flow-Noise 98.3 89.7 66.1 84.5 84.7 +29.1
(a) CALVIN
(b) ManiSkill
(c) MetaWorld
Figure 5: Comprehensive OOD evaluation results on CALVIN ABC-D, ManiSkill OOD, and MetaWorld ML45 benchmarks.
5.2 Main Results
In this section, we assess the in-distribution (ID) performance of $\pi_{\texttt{RL}}$ across various benchmarks, followed by an analysis of its out-of-distribution (OOD) generalization.
5.2.1 In-Distribution RL Training
As detailed in Tab.˜1, $\pi_{\texttt{RL}}$ yields substantial performance gains over SFT baselines across all evaluated benchmarks. Specifically, the $\pi_{0}$ model achieves a maximum average improvement of +29.2%, while the $\pi_{0.5}$ variant demonstrates a +31.0% increase in average success rate.
Specifically for LIBERO, we perform few-shot SFT on the $\pi_{0.5}$ model followed by RL optimization to achieve a 98.3% success rate, outperforming the 96.9% success rate of the full-dataset SFT baseline. These performance gains extend to other challenging environments, including ManiSkill with its 4,352 pick-and-place task combinations, MetaWorld featuring 50 distinct manipulation primitives, and CALVIN for long-horizon sequential tasks. See Appendix Appendix˜C for comprehensive experimental details.
5.2.2 Out-of-Distribution RL Evaluation
While previous experiments demonstrate that RL yields performance improvements, a critical question remains: does RL yield an enhanced policy, or simply overfit to the ID environment driven by provided rewards? In this section, we evaluate RL-finetuned policies in OOD scenarios, where the environment distribution or the task objective deviates from the ID training setup, to assess their generalization capabilities.
As illustrated in Fig.˜5, the performance gains achieved in the ID setting effectively transfer to OOD scenarios in ManiSkill and CALVIN, where the domain shift primarily stems from environmental variations. Conversely, for the OOD setting in MetaWorld, which involves distinct manipulation tasks, performance fluctuates without showing significant improvement. This finding suggests that the benefits of RL are primarily localized to action-level refinement rather than broader augmentation of cross-task generalization capabilities. See Appendix Appendix˜D for more details.
5.3 Ablation Study
Given that Flow-SDE achieves performance comparable to Flow-Noise while offering higher computational efficiency, we conduct our ablation studies with the Flow-SDE method. Specifically, we investigate the impact of critic designs, noise injection strategies, and MDP formulations, with additional results on RL algorithms and hyper-parameters provided in Appendix Appendix˜F.
(a) Eval
(b) Value Loss
(c) Explained Variance
Figure 6: Ablation on the critic design within Flow-SDE $\pi_{0}$ on the LIBERO-Long, indicating that the critic $V_{\text{vlm}}$ attached after the VLM exhibits superior performance. Furthermore, a four-layer MLP demonstrates stronger regression capability in $V_{\text{expert}}$.
(a) Eval
(b) Explained Variance
(c) Update Time
Figure 7: Ablation on the MDP formulation within Flow-SDE of $\pi_{0}$ on the LIBERO-Goal.
(a) Train
(b) Eval
5.3.1 Critic Design
Placement. We compare two critic placement strategies, one positioned after the action expert ($V_{\text{expert}}$) and the other after the VLM ($V_{\text{vlm}}$), with $\pi_{0}$ model on the LIBERO-Long task suite. As illustrated in Fig.˜6, we observe that $V_{\text{vlm}}$ exhibits slightly superior performance, lower value loss, and higher explained variance, despite not receiving the proprioceptive state as input. This advantage can be attributed to a key difference in their input: $V_{\text{vlm}}$ learns a direct mapping from observation to value, while $V_{\text{expert}}$ must contend with optimization challenges arising from coupled state and noisy action inputs.
Nevertheless, to align with the concept of the value function, we maintain the $V_{\text{expert}}$ architecture for the $\pi_{0}$, ensuring that state information is incorporated to estimate the value.
5.3.2 Flow Policy MDP
With the same fixed noise injection strategy, we evaluate the one-layer MDP of Flow-Noise with the two-layer MDP of Flow-SDE on the LIBERO-Goal, as shown in Fig.˜7.
Table 2: Ablation study of hyperparameters for Flow-SDE on the LIBERO-Spatial. Train refers to policy performance during the stochastic rollout phase, whereas Eval refers to performance during the deterministic evaluation phase.
Models Stage Hyperparameters
Noise Level Denoise Step Action Chunk
0.2 0.5 0.8 1 2 4 8 5 10 20
SFT Train 62.3 56.0 46.6 9.4 28.3 56.1 62.6 56.0 60.7 70.3
Eval 65.2 65.2 65.2 63.8 64.9 65.2 63.2 65.2 70.5 72.6
RL Train 59.5 93.5 95.3 73.8 90.8 93.5 84.3 93.5 93.3 87.5
Eval 73.1 94.5 98.1 88.5 97.0 94.5 86.7 94.5 95.5 89.2
(a) Train
(b) Eval
(c) Clipped Fraction
(a) Train
(b) Eval
(a) Eval
(b) Explained Variance
We observe that the one-layer formulation converges most rapidly, but the final success rates remain consistent across all three formulations. In terms of computational efficiency, the hybrid two-layer paradigm achieves a $2\times$ speedup over the standard approach due to its shorter effective MDP chain. Notably, the one-layer MDP yields no substantial wall-clock time advantage over the standard two-layer model, stemming from the requirement to recalculate full denoising trajectories for log likelihood estimation.
5.3.3 Stochasticity Injection
We compare fixed and learnable noise injection strategies using the Flow-SDE MDP formulation on the LIBERO-Long suite. To ensure a controlled comparison, we set the entropy bonus for the learnable noise to zero, aligning it with the fixed noise approach.
Specifically, we set the fixed noise to 0.5, and lower and upper bounds for the learnable noise log-variance to 0.08 and 0.16. As depicted in Fig.˜8, two noise strategies exhibit similar train performance at step 0, which indicates comparable noise magnitudes. Furthermore, the converged performance affirms the efficiency of both injection methods.
5.4 Hyper-Parameters
Building on the Flow-SDE with $\pi_{0}$ model, we investigate the influence of the noise level, denoise step, and action chunk on the LIBERO-Spatial benchmark. We denote the train stage as the phase where the policy generates stochastic actions for exploration, whereas the evaluation stage involves generating deterministic actions. The train and eval success rates for the SFT baseline and the RL fine-tuned model after 100 training steps are presented in Tab.˜2.
Noise Level. The noise level $a$ in the Flow-SDE is defined in Eq.˜8, which governs the noise injection magnitude during the denoising process. As shown in Tab.˜2, the evaluation performance of the SFT baseline remains identical across all noise levels due to its reliance on deterministic ODE sampling. Conversely, its training performance exhibits a clear degradation as noise increases. This is intuitive, as higher noise levels can distort the flow trajectory, leading to an inaccurate estimation of the marginal action distribution.
Extending this analysis to the RL fine-tuning stage highlights a critical trade-off: while lower noise levels mitigate exploration-induced performance degradation, they simultaneously constrain the capacity for RL refinement. This trade-off is empirically supported by Fig.˜9, which shows that training with minimal noise ($a=0.2$) leads to instability, characterized by a significantly higher clip fraction. We attribute this instability to the substantially larger gradient magnitudes associated with low-noise regimes.
Denoise Step. The denoise step $K$ defines the number of discretization steps for action generation and is critical for controlling the fidelity of the ODE-to-SDE transition in Eq.˜8. In Tab.˜2, we observe that while all configurations start with similar eval performance, the train success rate plummets at $K=1$, indicating a significant ODE-to-SDE discretization error.
However, a larger $K$ is not unequivocally optimal. As illustrated in Fig.˜10, increasing $K$ presents a distinct trade-off: while it enhances rollout performance, it simultaneously increases training complexity and computational overhead due to the extended sequence of denoising steps.
Action chunk. The action chunk refers to the number of consecutive actions the policy executes within a single observation. We ablate the action chunk size across 5, 10, and 20, with results visualized in Fig.˜11.
Although a larger chunk size yields marginal performance gains, it inherently reduces the frequency of policy-environment interactions and obscures precise reward credit assignment. These constraints lead to less reliable advantage estimation, as evidenced by the diminished explained variance. Consequently, while an increased chunk size may offer a superior SFT baseline, it paradoxically limits the ceiling for subsequent RL-driven refinement.
6 Conclusion
We introduce $\pi_{\texttt{RL}}$, a framework that enables flow-based VLAs, $\pi_{0}$ and $\pi_{0.5}$, to be fine-tuned with online RL algorithms. We tackle the fundamental challenge of intractable log-likelihoods in flow matching with Flow-Noise and Flow-SDE solutions. Our extensive experiments on the challenging benchmarks demonstrated that $\pi_{\texttt{RL}}$ achieves significant performance improvements over SFT baselines.
Limitation. Due to the low sample efficiency of online RL, our framework currently relies on sim-to-real deployment. We aim to develop more efficient algorithms to enable real-world RL training in the future.
Appendix A Appendix
This appendix provides additional technical details and experimental results for $\pi_{\texttt{RL}}$. The content is organized as:
  • Appendix B: Experimental Setup.
  • Appendix C: In-Distribution RL Training.
  • Appendix D: Out-of-Distribution RL Evaluation.
  • Appendix E: Case Studies: Single-Task RL Training.
  • Appendix F: Ablation Details.
  • Appendix G: Insights from Large-Scale RL Training.
  • Appendix H: RL for GR00T N1.5.
  • Appendix I: Limitations and Future Work.
  • Appendix J: Training Hyperparameters.
Appendix B Experimental Setup
B.1 Benchmarks.
To rigorously assess the performance and generalization of our framework, we conduct evaluations across four diverse benchmarks, with diverse emphases on robotic capability:
  • LIBERO focuses on compositional variations and knowledge transfer in tasks. By evaluating across four task suites: Spatial, Object, Goal, and Long, it probes the model’s ability to transfer base skills to variations in object arrangements and long-horizon tasks.
  • ManiSkill emphasizes perceptual and execution robustness under massive environmental diversity. We follow the setup of RL4VLA (Liu et al., 2025b) and built a benchmark with 4,352 unique pick-and-place combinations, which challenges the policy to maintain physical interactions across a vast distribution of objects and receptacles.
  • CALVIN evaluates long-horizon sequential reasoning and vision-language grounding. It evaluates the model’s capacity to execute chains of five random subtasks in a persistent environment, requiring accurate alignment with linguistic instructions.
  • MetaWorld measures skill breadth and multi-task versatility. It requires a single policy to master 50 semantically distinct manipulation primitives, ranging from simple reaching to complex tool usage.
B.2 Implementation Details.
@@ Appendix B Experimental Setup
Given that pre-trained models often struggle to generalize to task-specific benchmarks, we initiate our process with SFT on expert demonstrations. For the SFT stage, we fine-tune the entire 3.3B model following the official setting. In the subsequent RL stage, we freeze the VLM parameters and exclusively fine-tune the 300M action expert model, driven by GPU memory efficiency and the findings from RL4VLA that RL contributes more significantly to action generalization. We build the whole framework upon the RLinf (Yu et al., 2025) codebase, where we adopt a shared, co-located GPU allocation strategy that places the environment, rollout model, and actor model on the same GPU and executes them serially.
@@
For the model configurations, we adhere to the official setting provided by openpi (Black et al., 2024; Intelligence et al., 2025). In these settings, $\pi_{0}$ utilizes image, language, and proprioceptive states as input, whereas $\pi_{0.5}$ notably omits state information [^2] https://github.com/Physical-Intelligence/openpi/issues/687. Our experiments are conducted on 8 NVIDIA H100 80GB GPUs, and detailed training hyperparameters are available in the Appendix Appendix˜J.
@@
Table 4: Evaluation results on the ManiSkill benchmark, with more specific OOD results depicted in Tab.˜7.
Model IND OOD
Avg. $\Delta$ Vision Semantic Execution Avg. $\Delta$
$\pi_{0}$ SFT 38.4 32.6 8.4 13.2 18.1
Flow-SDE 78.8 +40.4 61.1 25.4 31.5 39.3 +21.2
Flow-Noise 77.8 +39.4 63.4 23.1 24.2 36.9 +18.8
$\pi_{0.5}$ SFT 40.1 40.2 16.6 22.4 26.4
Flow-SDE 90.9 +50.8 68.0 34.5 45.4 49.3 +22.9
Flow-Noise 89.7 +49.6 69.9 35.5 54.9 53.4 +27.0
Appendix C In-Distribution RL Training
In this section, we evaluate the effectiveness of our framework across various benchmarks to demonstrate its robustness and superior performance in ID settings.
C.1 LIBERO Benchmark